OpenStudy (anonymous):

Which is the equation to the line passing through points (2, -7) and (5, -1)? A) y = 2x - 8 B) y = - 2x + 9 C) y = 2x - 11 D) y = - 2x + 6

OpenStudy (anonymous):

y2-y1/x2-x1=m substitute the point (-1-(-7)/(5-2)=m 66/3=m 2=m then to find equation y-y1=m(x-x1) y-(-7)=2(x-2) y+7=2x-4 y=2x-11 answer c